# Two-Photon Absorption in Silicon Using Real Density Matrix Approach

### v1

@@ II.4 Nonlinear absorption coefficients
$\alpha_{2,em}(\omega_{a},\omega_{b})$
$=-\frac{4E_{g}(n_{ph}+1)}{\epsilon_{0}\hbar\,n_{a}\,n_{b}\,c^{2}}M_{0}^{4}\left(\frac{2\mu_{\parallel H}}{\hbar^{2}}\right)^{2}\frac{1}{4\pi r_{0}}\frac{T_{1}}{\hbar}$
×$\times x_{a}\left(\frac{\sin(\tilde{\kappa}_{a-,em}r_{0})}{\tilde{\kappa}_{a-,em}r_{0}}\right)^{2}[\tilde{\kappa}_{b-,em}r_{0}+\tilde{\kappa}_{b-,abs}r_{0}]a_{H}^{*2}$
×$\times\frac{A^{\prime\prime}_{e}+A^{\prime\prime}_{h}}{a_{H}^{*2}},$ (60)
$\alpha_{{2c,abs}}(\omega_{a},\omega_{b})$
$=-\frac{4E_{g}\,n_{ph}}{\epsilon_{0}\hbar\,n_{a}\,n_{b}\,c^{2}}M_{0}^{4}\left(\frac{2\mu_{\parallel H}}{\hbar^{2}}\right)^{2}\frac{1}{4\pi r_{0}}\frac{T_{1}}{\hbar}$
×$\times x_{a}\left(\frac{\sin(\tilde{\kappa}_{a-,abs}r_{0})}{\tilde{\kappa}_{a-,abs}r_{0}}\right)^{2}[\tilde{\kappa}_{b-,em}r_{0}+\tilde{\kappa}_{b-,abs}r_{0}]a_{H}^{*2}$
×$\times\frac{A^{\prime\prime}_{e}+A^{\prime\prime}_{h}}{a_{H}^{*2}}.$ (61)
The solutions obtained in the two regimes of discrete and continuous states are smoothly connected with the use of hyperbolic tangent function. Using the above formula, we have calculated the 2PA coefficient $\alpha_{2}(\omega,\omega_{b})$ as a function of the energies $\hbar\omega_{a}+\hbar\omega_{b}$. The band parameters used in the calculations are listed in Tables 1 - 3.
Table 2: Anisotropy parameters for Si, excitonic energies calculated from Eq. (II.2)
Parameter
      Value
Unit
$\eta_{00v}=\eta_{v}$
$\gamma_{aH}(4.2\hbox{K})$
      0.43
1.13
$|E_{1H}|\,(4.2\hbox{K})$
      17.38
meV
$\gamma_{aH}(300\hbox{K})$
      0.29
1.18
$|E_{1H}|\,(300\hbox{K})$
      21.75
meV
$\gamma_{aL}(4.2\hbox{K})$
      0.64
1.07
$|E_{1L}|\,(4.2\hbox{K})$
      9.94
meV
$\gamma_{aL}(300\hbox{K})$
      0.62
1.078
$|E_{1L}|\,(300\hbox{K})$
      9.70
meV
Table 3: Thermal electron and hole lengths for Si, and expressions $\varphi_{1H}(0)A_{1H},B_{1H}\;\varphi_{2H}(0)A_{2H},B_{2H}$
Quantity
Value 4.2K
Value 300 K
Reference
$\overline{\lambda}_{eH}$
5.09
0.675
$\tilde{\lambda}_{hH}$
4.70
0.42
$\varphi_{1H}(0)A_{1H}$
$4\times 1.64$
$4\times 0.23$
$\varphi_{1H}(0)B_{1H}$
$4\times 1.57$
$4\times 0.155$
@@ III Experimental
We perform optical experiments to determine nonlinear absorption coefficients of silicon to supplement existing data in the literature, spanning a wider range of $x=\frac{\hbar(\omega_{a}+\omega_{b})}{2E_{g}}$ values. For this purpose, we perform cross correlation experiments, using optical pulses derived from 1-kHz amplified femtosecond laser system (Spitfire Ace, Spectra Physics) system. The laser seeds two optical parametric amplifiers (OPA, TOPAS-Prime, Light Conversion), where one OPA is used as a source of near-infrared (NIR) probe radiation between 1150 nm - 1350 nm (1.08 eV - 0.91 eV), producing pulses in the range of 100-150 fs. The signal and idler pulses from the second OPA system are used to generate MIR pump pulses through the process of difference frequency generation (DFG) in the 2480 nm to 4651 nm (0.5 eV - 0.27 eV) range, producing 175-265 fs pulses. The MIR pump is modulated using an optical chopper at a frequency of 500 Hz that is synchronized to the laser output (MC2000B, Thorlabs).
We use a 280 $\mu$m thick [100] silicon window (99.999% purity, University Wafer) as the sample target. The MIR and NIR pulses are focused in a non-collinear arrangement on the Si target, using normal incidence for the pump beam and a $\sim 20$ degree incidence angle for the probe beam. Temporal overlap is controlled through an automated translation stage (GTS150, Newport) in the probe arm, producing a cross-correlation of pump induced probe absorption via NTA. The remaining probe is then attenuated by a 3 OD neutral density filter and detected using a home-built InGaAs photodiode. The modulated change induced by the MIR pump is analyzed by a lock-in amplifier (SR860, Stanford Research Systems). The resulting cross-correlation is then used to extract a nonlinear absorption coefficient $\alpha_{2}$ as described by Negres et al. [34]
@@
Figure 1: Exemplary pump probe data and corresponding nonlinear transmittance fit. The photon energies used here are 0.91 eV and 0.376 eV.
@@ IV Results and Discussion
The derived form of $\alpha_{2}$ that follows from the RDMA approach contains contributions from the two regimes of exciton production, namely the contributions from discrete states and continuum states, see equation (57). Both regimes rely on dipole allowed transitions while considering either the absorption or emission. The discrete regime is visually depicted in Figure 2, which highlights the two-dimensional dispersion of $\alpha_{2}$. The analysis recognizes the presence of a single discrete excitonic state below the band edge, corresponding to the creation of a bound state through absorption of a phonon, $\hbar\Omega_{abs}$, which may provide considerable enhancement to the 2PA process. This is manifested by the presence of resonant denominators for the discrete regime, which can be written in simplified form from equation (II.4) as
$\alpha_{2d}(\omega_{a},\omega_{b})\propto\frac{1}{(\hbar\Omega_{abs}/E_{g}-\hbar\omega_{b}/E_{g})^{2}+\gamma_{2}^{2}}\times$
@@
Figure 2: Scaling behavior of the two photon absorption coefficient $\alpha_{2}$ as a function of normalized photon energy while both photons are below the gap.
If the the energy of the incident photons exceeds the bandgap energy, then the two-photon absorption process proceeds entirely via continuum states. This can be accessed by either phonon absorption or emission, and the nonlinear absorption coefficient has the form as in equation (59). Beyond the band edge, the continuum causes $\alpha_{2c}$ to further increase up to the point where the energy of at least one of the photons has enough energy to reach the direct gap of silicon at $E_{g}=3.43~{}\rm{eV}$. This increase is due to the oscillatory behavior of the linear coherent amplitude of the exciton density $Y^{(1)}$, which contributes to the induced polarization of the medium in order for two-photon absorption to occur.
@@
We next use the model to describe the energy scaling of the NTA coefficient as measured in Ref [23], which encompasses a normalized photon energy range of $x=\hbar(\omega_{a}+\omega_{b})/2E_{g}=1.1-2$. In addition, we apply the model to fit a new data set that we have collected for a non-degenerate energy ratio in the $\hbar\omega_{a}/\hbar\omega_{b}=2-4$ range, corresponding to pump photon energies of 0.294 eV–0.5 eV and probe energies of 0.91 eV–1.08 eV. Both data sets, shown in Figures 4 and 5 , correspond to a normalized equivalent energy range of $x=0.55-0.8$, but have different levels of non-degeneracy, thus complementing each other. The theoretical predictions for $\alpha_{2}$ are in good agreement with both experimental data sets. The most notable feature in the normalized photon energy curve is the appearance of a peak, which the model attributes to the presence of the bound exciton resonance. The position of this peak shifts in accordance with the energy tuning of the lower energy pump photon, and its magnitude is largely dictated by the damping terms in the resonant denominators of $\alpha_{2d}$. However, this resonant behavior is significantly different from what is reported in [23], where the theoretical predictions made anticipate a quasi-linear scaling of $\alpha_{2}$ as a function of the normalized photon energy. Future work may include performing NTA experiments to verify the behavior of $\alpha_{2}$ and the influence of continuum states above the band edge, such as can be obtained with a two-color Z-scan.
Figure 4: Scaling of the non-degenerate two-photon absorption coefficient $\alpha_{2}$ as a function of the normalized equivalent energy. Experimental data points are obtained from Ref [23] and thick solid lines represent the results from the RDMA analysis. Different colors correspond to different non-degeneracy energy ratios $\hbar\omega_{a}/\hbar\omega_{b}$.
This work is a significant departure from the conventional understanding of the two-photon absorption process in semiconductors. Previous methods have described the 2PA absorption process in terms of transition rate matrix elements based on electron band states, where significantly nondegenerate photons become resonant with the interband transition and an intraband transition simultaneously. Within the framework of the RDMA, the defining feature of the absorption process is the ability to produce either bound or free excitons within the material. Tracking the formation of excitons allows the RDMA to yield general expressions for semiconducting materials with low binding energies, i.e. only the lowest discrete state needs to be considered. These results are analytical in that there is no parametrization done to provide free variables for fitting. The only variability results from loosely determined material parameters, namely the relaxation times $T_{1}$ and $T_{2}$.
@@
Figure 5: Scaling of the non-degenerate two-photon absorption coefficient $\alpha_{2}$ as a function of the normalized equivalent energy. Data points indicate experimental results and thick solid lines represent the results from the RDMA analysis. Different colors correspond to different non-degeneracy energy ratios $\hbar\omega_{a}/\hbar\omega_{b}$.
@@ V Conclusion
We employed the RDMA approach to elucidate two-photon absorption (2PA) in indirect gap semiconductors, employing silicon as the representative material. Our methodology, in contrast to existing models for 2PA in semiconductors, uniquely incorporates a detailed description of DTA (direct two-photon absorption) and NTA (non-direct two-photon absorption) through the production of excitons, a physical insight often overlooked in prior analyses. The approach allows the computation of the non-degenerate two-photon absorption coefficient, capturing the influence of both bound and free excitonic states within the material. In addition, the RDMA enables the inclusion of the effects of phonon-assisted transitions and material anisotropy. According to our model, the two-photon absorption process in silicon intensifies as the energy of individual incident photons, while below the gap, approaches the single bound excitonic state. This prediction aligns with recently obtained experimental NTA absorption data, casting new light on the interpretation of published DTA data in bulk silicon. Above the gap, the 2PA process primarily occurs through the continuum of free states, providing a satisfactory description of the DTA data across a wide photon energy range ($\hbar\omega/E_{g}=0.5-1.3$). Last, but not least, the analysis outlined using the RDMA can be generalized and readily applied to other indirect gap materials by adjusting the necessary material values. Furthermore, the model can be adapted to describe direct transitions by excluding the involvement of phonon modes to complete the transition.

### v2

@@ II.6 Nonlinear absorption coefficients
$\alpha_{2cH,em}(\omega_{a},\omega_{b})$
$=-\frac{4E_{g}(n_{ph}+1)}{\epsilon_{0}\hbar\,n_{a}\,n_{b}\,c^{2}}M_{0H}^{4}\left(\frac{2\mu_{\parallel H}}{\hbar^{2}}\right)^{2}\frac{1}{4\pi r_{0H}}\frac{T_{1}}{\hbar}$
×$\times x_{a}\left(\frac{\sin(\tilde{\kappa}_{aH-,em}r_{0H})}{\tilde{\kappa}_{aH-,em}r_{0H}}\right)^{2}[\tilde{\kappa}_{bH-,em}r_{0H}+\tilde{\kappa}_{bH-,abs}r_{0H}]$
×$\times\left({{\mathcal{A}}^{\prime\prime}_{eH,em}+{\mathcal{B}}^{\prime\prime}_{hH,em}}\right),$ (105)
and
$\alpha_{{2cH,abs}}(\omega_{a},\omega_{b})$
$=-\frac{4E_{g}\,n_{ph}}{\epsilon_{0}\hbar\,n_{a}\,n_{b}\,c^{2}}M_{0H}^{4}\left(\frac{2\mu_{\parallel H}}{\hbar^{2}}\right)^{2}\frac{1}{4\pi r_{0H}}\frac{T_{1}}{\hbar}$
×$\times x_{a}\left(\frac{\sin(\tilde{\kappa}_{aH-,abs}r_{0H})}{\tilde{\kappa}_{aH-,abs}r_{0H}}\right)^{2}[\tilde{\kappa}_{bH-,em}r_{0H}+\tilde{\kappa}_{bH-,abs}r_{0H}]$
×$\times\left({{\mathcal{A}}^{\prime\prime}_{eH,abs}+{\mathcal{B}}^{\prime\prime}_{hH,abs}}\right).$ (106)
The solutions obtained in the two regimes of discrete and continuous states are smoothly connected via the use of hyperbolic tangent functions. Using the above formula, we have calculated the 2PA coefficient $\alpha_{2}(\omega_{a},\omega_{b})$ as a function of the energies $\hbar\omega_{a}+\hbar\omega_{b}$. The band parameters used in the calculations are listed in Tables 1 - 3.
The masses are calculated from Luttinger parameters
$m_{hzH}$ $=$ $\frac{m_{0}}{\gamma_{1}-2\gamma_{2}},$
$m_{h\parallel H}$ $=$ $\frac{m_{0}}{\gamma_{1}-2\gamma_{3}},$
$m_{hzL}$ $=$ $\frac{m_{0}}{\gamma_{1}+2\gamma_{2}},$ (107)
$m_{h\parallel L}$ $=$ $\frac{m_{0}}{\gamma_{1}+2\gamma_{3}}.$
With the reduced electron and hole masses, the thermal lengths are given by
$\overline{\lambda}_{th,e}=\left(\frac{\hbar^{2}}{m_{e\parallel}k_{B}{\mathcal{T}}}\right)^{1/2}=\left(2\times\frac{\hbar^{2}}{2\mu_{\parallel H}}\times\frac{\mu_{\parallel H}}{m_{e\parallel}}\frac{1}{k_{B}{\mathcal{T}}}\right)^{1/2}$
$=\left(2\times\frac{\mu_{\parallel H}}{m_{e\parallel}}\frac{R^{*}_{H}}{k_{B}{\mathcal{T}}}\right)^{1/2},$
$\overline{\lambda}_{th,hH}=\left(\frac{\hbar^{2}}{m_{h\parallel}k_{B}{\mathcal{T}}}\right)^{1/2}$
$k_{B}T=8.617\times 10^{-2}\;\frac{\hbox{meV}}{\hbox{K}}.$ (108)
The calculated values are summarized in Table 3.
Table 3: Thermal electron and hole lengths for Si, and expressions $\varphi_{1H}(0)A_{1H}\;,\varphi_{1H}(0)B_{1H}$
Quantity Value 4.2K Value 300 K Reference
$\overline{\lambda}_{th,e\parallel}$ 8.08 1.05 Eq. (II.6)
$\overline{\lambda}_{th,ez}$ 3.68 0.45 Eq. (II.6)
$\overline{\lambda}_{th,h\parallel H}$ 4.15 0.42 Eq. (II.6)
$\overline{\lambda}_{th,hzH}$ 6.66 0.67 Eq. (II.6)
$\varphi_{1H}(0)A_{1H}$ $4\times 1.52$ $4\times 0.173$
$\varphi_{1H}(0)B_{1H}$ $4\times 1.34$ $4\times 0.06$
@@ III Experimental
We perform optical experiments to determine nonlinear absorption coefficients of silicon to supplement existing data in the literature, spanning a wider range of $x=\frac{\hbar(\omega_{a}+\omega_{b})}{2E_{g}}$ values. For this purpose, we perform cross correlation experiments, using optical pulses derived from a 1-kHz amplified femtosecond laser system (Spitfire Ace, Spectra Physics). The laser seeds two optical parametric amplifiers (OPA, TOPAS-Prime, Light Conversion), where one OPA is used as a source of near-infrared (NIR) probe radiation between 1150 nm - 1350 nm (1.08 eV - 0.91 eV), producing pulses in the range of 100-150 fs. The signal and idler pulses from the second OPA system are used to generate MIR pump pulses through the process of difference frequency generation (DFG) in the 2480 nm to 4651 nm (0.5 eV - 0.27 eV) range, producing 175-265 fs pulses. The MIR pump is modulated using an optical chopper at a frequency of 500 Hz that is synchronized to the laser output (MC2000B, Thorlabs).
We use a 280 $\mu$m thick [100] silicon window (99.999% purity, University Wafer) as the sample target. The MIR and NIR pulses are focused in a non-collinear arrangement on the Si target, using normal incidence for the pump beam and a $\sim 20$ degree incidence angle for the probe beam. Temporal overlap is controlled through an automated translation stage (GTS150, Newport) in the probe arm, producing a cross-correlation of pump induced probe absorption via NTA. The remaining probe is attenuated by an OD = 3 neutral density filter and detected using a home-built InGaAs photodiode. The modulated change induced by the MIR pump is analyzed by a lock-in amplifier (SR860, Stanford Research Systems). The resulting cross-correlation is then used to extract the nonlinear absorption coefficient $\alpha_{2}$ as described by Negres et al. [47]
@@
Figure 1: Exemplary pump-probe data and corresponding nonlinear transmittance fit. The photon energies used here are 0.91 eV and 0.376 eV.
@@ IV Results and Discussion
The derived form of $\alpha_{2H}$ that follows from the RDMA contains contributions from the two regimes of exciton production, namely the contributions from discrete states and continuum states, see equation (102). The discrete regime is visually depicted in Figure 2, which highlights the two-dimensional dispersion of $\alpha_{2}$. The analysis recognizes the presence of a single discrete excitonic state below the band edge, corresponding to the creation of a bound state through absorption of a phonon, $\hbar\Omega_{abs}$, which may provide considerable enhancement to the 2PA process. This is manifested by the presence of resonant denominators for the discrete regime, which can be written in simplified form from equation (II.6) as
$\alpha_{2dH}(\omega_{a},\omega_{b})\propto\frac{1}{(\hbar\Omega_{abs}/E_{g}-\hbar\omega_{b}/E_{g})^{2}+\gamma_{2}^{2}}\times$
@@
Figure 2: Scaling behavior of the two photon absorption coefficient $\alpha_{2}$ (Eq. 103) as a function of normalized photon energy while both photons are below the gap.
If the the energy of the incident photons exceeds the bandgap energy, then the two-photon absorption process proceeds entirely via continuum states. This can be accessed by either phonon absorption or emission, and the nonlinear absorption coefficient has the form as in equation (104). Beyond the band edge, the continuum causes $\alpha_{2cH}$ to further increase up to the point where the energy of at least one of the photons has enough energy to reach the direct gap of silicon at $E_{g}=3.43~{}\rm{eV}$. This increase is due to the oscillatory behavior of the linear coherent amplitude of the exciton density $Y^{(1)}$, which contributes to the induced polarization of the medium in order for two-photon absorption to occur.
@@
We next use the model to describe the energy scaling of the NTA coefficient as measured in Ref [23], which encompasses photons with energy ratio from 1.6 to 2.08 (Figure 4). This compliments our study using pump photon energies of 0.294 eV–0.5 eV and probe energies of 0.91 eV–1.08 eV, corresponding to photon energy ratio from 1.82 to 3.67 (Figure 5). The theoretical predictions for $\alpha_{2}$ are in good agreement with both experimental data sets. The most notable feature in the normalized photon energy curve is the appearance of a peak, which the model attributes to the presence of the bound exciton resonance. The position of this peak shifts in accordance with the energy tuning of the lower energy pump photon, and its magnitude is largely dictated by the damping terms in the resonant denominators of $\alpha_{2dH}$. However, this resonant behavior is significantly different from what is reported in [23], where the theoretical predictions made anticipate a quasi-linear scaling of $\alpha_{2}$ as a function of the normalized photon energy. Future work may include NTA experiments to verify the behavior of $\alpha_{2}$ and the influence of continuum states above the band edge, such as can be obtained with a two-color Z-scan.
Figure 4: Scaling of the non-degenerate two-photon absorption coefficient $\alpha_{2}$ as a function of the normalized equivalent energy. Experimental data points are obtained from Ref [23] and thick solid lines represent the results from the RDMA analysis. Different colors correspond to different photon energies $\hbar\omega_{a}$.
This work is a significant departure from the conventional understanding of the two-photon absorption process in semiconductors. Previous methods have described the 2PA absorption process in terms of transition rate matrix elements based on electron band states, where significantly non-degenerate photons become resonant with the interband transition and an intraband transition simultaneously. Within the framework of the RDMA, the defining feature of the absorption process is the ability to produce either bound or free excitons within the material. Tracking the formation of excitons allows the RDMA to yield general expressions for semiconducting materials with low binding energies, i.e. only the lowest discrete state needs to be considered. These results are analytical in that there is no parametrization done to provide free variables for fitting. The only variability results from loosely determined material parameters, namely the relaxation times $T_{1}$ and $T_{2}$.
@@
Figure 5: Scaling of the non-degenerate two-photon absorption coefficient $\alpha_{2}$ as a function of the normalized equivalent energy. Data points indicate experimental results obtained in current study and thick solid lines represent the results from the RDMA analysis. Different colors correspond to different photon energies $\hbar\omega_{a}$.
@@ V Conclusion
We have employed the RDMA approach to elucidate two-photon absorption (2PA) in indirect gap semiconductors, using silicon as the representative material. Our methodology, in contrast to existing models for 2PA in semiconductors, uniquely incorporates a detailed description of DTA (degenerate two-photon absorption) and NTA (non-degenerate two-photon absorption) through the production of excitons, a physical insight often overlooked in prior analyses. The approach allows the computation of the non-degenerate two-photon absorption coefficient, capturing the influence of both bound and free excitonic states within the material. In addition, the RDMA enables the inclusion of the effects of phonon-assisted transitions and material anisotropy. According to our model, the two-photon absorption process in silicon intensifies as the energy of individual incident photons, while below the gap, approaches the single bound excitonic state. This prediction aligns with recently obtained experimental NTA absorption data, shedding new light on the interpretation of published DTA data in bulk silicon. Above the gap, the 2PA process primarily occurs through the continuum of free states, providing a satisfactory description of the DTA data across a wide range of photon energies ($\hbar\omega/E_{g}=0.5-1.3$). Last, but not least, the analysis outlined using the RDMA can be generalized and readily applied to other indirect gap materials by adjusting the necessary material values. Furthermore, the model can be adapted to describe direct transitions by excluding the involvement of phonon modes to complete the transition.
@@ Acknowledgements.
Data Availability Statement
The data that support the findings of this study are available from the corresponding author upon reasonable request.